Find the equation for the line that passes through the point (−5, 5), and that is perpendicular to the line with the equation −9/4x − 3y = −18
Answer by Cromlix(4381)

Hi there,
−9/4x − 3y = −18
Rearrange to y = mx + c form
-3y = 9/4x - 18
Multiply throughout by -1
3y = -9/4x + 18
Divide throughout by 3
y = -9/12x + 6
y = -3/4x + 6
Lines that are perpendicular to
one another have gradients that
multiply together to give -1
m1 x m2 = -1
-3/4 x m2 = -1
m2 = 4/3
Using line equation:
y - b = m(x - a)
Where, m = 4/3 and (a,b) = (-5,5)
y - 5 = 4/3(x - (-5))
y - 5 = 4/3(x + 5)
y - 5 = 4/3x + 20/3
y = 4/3x + 20/3 + 15/3 (5)
y = 4/3x + 35/3
or
3y = 4x + 35.
Hope this helps :-)